If a train travels M miles in H hours and 45 minutes, what is the average speed in miles per hour?
Answer by stanbon(75887)

If a train travels M miles in H hours and 45 minutes, what is the average speed in miles per hour?
----
distance = M miles
time = (3/4) hr
rate = d/t = M/(3/4) = (4/3)M mph
================================================================
Cheers,
Stan H.